Check whether 𝑇 ∶ ℝ2 → ℝ2




, defined by 𝑇 (𝑥, 𝑦) = (−𝑦, 𝑥) is a linear transformation.

1
Expert's answer
2022-05-20T07:54:40-0400

A linear transformation (or a linear map) is a function "\ud835\udc47 \u2236 \u211d^2 \u2192 \u211d^2" that satisfies the following properties:


"T(a+b)=T(a)+T(b)"

"T(\alpha a)=\alpha T(a)"

for any vectors  "a, b\in \R^2" and any scalar "\alpha\in \R."

Let "a=(x_1, y_1), b=(x_2, y_2)." Then


"T(a)=(-y_1, x_1)"

"T(b)=(-y_2, x_2)"

"T(a+b)=(-(y_1+y_2), x_1+x_2)"

"=(-y_1, x_1)+(-y_2, x_2)=T(a)+T(b), True"

"T(\alpha a)=(-\alpha y_1,\alpha x_1)=\alpha(-y_1, x_1)"

"=\alpha T(a), True"

Therefore a linear transformation "\ud835\udc47 \u2236 \u211d^2 \u2192 \u211d^2"defined by "T(x, y)=(-y,x)" is a linear transformation.